OpenStudy (anonymous):

Eliminate the parameter. x = 5t, y = t + 8

OpenStudy (anonymous):

I will give medal

OpenStudy (anonymous):

\[ 5 x =t \implies t =\frac x 5 \] Can you finish it?

OpenStudy (anonymous):

How much is y=t + 8 in terms of x?

OpenStudy (anonymous):

\[ y= t+8=\frac x 5 +8 \]

OpenStudy (anonymous):

dang u good at math